OpenStudy (anonymous):

Solve for x. 10 - |x + 2| = 12

OpenStudy (anonymous):

no solution

OpenStudy (amorfide):

there is no solution?

OpenStudy (amorfide):

if you move the constants to one side so subtract 10 from both sides -|x+2|=2 you must get -2 for |x+2| for this to be true but |x| is always positive

OpenStudy (anonymous):

x=> -2 -> 10 -(x+2) = 12 10 -x -2 = 12 8 -x =12 x = -20 but since we are in a part which x=> -2 it cant be .. x < -2 10 - (-(x+2)) = 12 10 + x + 2 = 12 12 + x = 12 x = 0 but this cant be as well for the same reason .. no solution

OpenStudy (anonymous):

okay thanks!